# SwiftSeg: Efficient Training-Free Open-Vocabulary Segmentation via Hierarchical Attention Refinement Method

### v2

@@ 5.2 Conclusion
In this paper, we introduced SwiftSeg, a novel training-free open-vocabulary semantic segmentation (OVSS) method that leverages diffusion models for efficient and accurate segmentation. Our method combines (1+1)-step DDIM inversion, a dual-prompt mechanism, and a multi-resolution attention fusion strategy. Unlike prior approaches, SwiftSeg generates segmentation maps for all candidate classes in a single inference run, significantly reducing computational overhead. Additionally, we incorporate Test-Time Flipping (TTF) to improve prediction robustness. Extensive experiments on PASCAL VOC, PASCAL Context, and COCO Object benchmarks demonstrate that SwiftSeg achieves state-of-the-art performance among training-free diffusion-based methods, with the highest average mIoU and achive inference time under one seconds. These results highlight the practicality and scalability of SwiftSeg for real-world OVSS applications.
Building upon SwiftSeg, future research can explore several directions to further enhance segmentation quality and flexibility. One promising avenue is the development of adaptive candidate class generation mechanisms, which leverage contextual cues to dynamically refine class sets and reduce dependency on the auxiliary candidate class generation process. Additionally, extending SwiftSeg to support instance-level segmentation would enable more fine-grained visual understanding. With its high adaptability, accurate segmentation performance, and low segmentation latency, SwiftSeg provides a practical and deployable solution for automatic segmentation in real-world applications.

### v3

@@ 5.2 Conclusion
In this paper, we introduced FA-Seg, a novel training-free open-vocabulary semantic segmentation (OVSS) method that leverages diffusion models for efficient and accurate segmentation. Our method combines (1+1)-step DDIM inversion, a dual-prompt mechanism, and a multi-resolution attention fusion strategy. Unlike prior approaches, FA-Seg generates segmentation maps for all candidate classes in a single inference run, significantly reducing computational overhead. Additionally, we incorporate Test-Time Flipping (TTF) to improve prediction robustness. Extensive experiments on PASCAL VOC, PASCAL Context, and COCO Object benchmarks demonstrate that FA-Seg achieves state-of-the-art performance among training-free diffusion-based methods, with the highest average mIoU and achive inference time under one seconds. These results highlight the practicality and scalability of FA-Seg for real-world OVSS applications.
Building upon FA-Seg, future research can explore several directions to further enhance segmentation quality and flexibility. One promising avenue is the development of adaptive candidate class generation mechanisms, which leverage contextual cues to dynamically refine class sets and reduce dependency on the auxiliary candidate class generation process. Additionally, extending FA-Seg to support instance-level segmentation would enable more fine-grained visual understanding. With its high adaptability, accurate segmentation performance, and low segmentation latency, FA-Seg provides a practical and deployable solution for automatic segmentation in real-world applications.
Figure 10: Visualization of some failure cases by FA-Seg
6 Acknowledgement
This research was supported by The VNUHCM-University of Information Technology’s Scientific Research Support Fund.
[…]